OpenStudy (anonymous):

Marta is making a scale drawing of her bedroom. Her rectangular room is 12.5 feet wide and 15 feet long. On the scale drawing, the width of her room is 5 inches. What is the length on her drawing? Please explain! I am not here for just answers. I have tried cross multiplying! I get 3.5 The only answers I have to choose from is 6/5.5/4.5/10.

OpenStudy (stamp):

\[W_{actual}=12.5\ ft\]\[W_{scale}=5\ in\]\[L_{actual}=15\ ft\]\[L_{scale}=x\ in\]Set up a proportion of scale/actual \[\frac{W_{scale}}{W_{actual}}=\frac{L_{scale}}{L_{actual}}\]\[\frac{5}{12.5}=\frac{x}{15}\]Solve for x.

OpenStudy (anonymous):

I tried that! I still get 3.5!

OpenStudy (stamp):

As far as test taking goes, if the actual length is > than actual width, the scale length should be > scale width. So if scale width = 5 we can rule out scale length = 4.5.

OpenStudy (stamp):

\[5(15)=\frac{25}{2}x\]\[75=\frac{25}{2}x\]\[75(2)=25x\]\[150=25x\]\[x=\frac{150}{25}\]\[x=6\ in\]